# Impulsive mixing of stellar populations in dwarf spheroidal galaxies

### v1

@@ Abstract
We study the response of mono-energetic stellar populations with initially isotropic kinematics to impulsive and adiabatic changes to an underlying dark matter potential. Half-light radii expand and velocity dispersions decrease as enclosed dark matter is removed. The details of this expansion and cooling depend on the time scale on which the underlying potential changes. In the adiabatic regime, the product of half-light radius and average velocity dispersion is conserved. We show that the stellar populations maintain centrally isotropic kinematics throughout their adiabatic evolution, and their densities can be approximated by a family of analytical radial profiles. Metallicity gradients within the galaxy flatten as dark matter is slowly removed. In the case of strong impulsive perturbations, stellar populations develop power-law-like density tails with radially biased kinematics. We show that the distribution of stellar binding energies within the dark matter halo substantially widens after an impulsive perturbation, no matter the sign of the perturbation. This allows initially energetically separated stellar populations to mix, to the extent that previously chemo-dynamically distinct populations may masquerade as a single population with large metallicity and energy spread. Finally, we show that in response to an impulsive perturbation, stellar populations that are deeply embedded in cored dark matter halos undergo a series of damped oscillations before reaching a virialised equilibrium state, driven by inefficient phase mixing in the harmonic potentials of cored halos. This slow return to equilibrium adds substantial systematic uncertainty to dynamical masses estimated from Jeans modeling or the virial theorem.
Cold dark matter (265); Dwarf spheroidal galaxies (420); Galaxy dynamics (591); N-body simulations (1083); Perturbation methods (1215)
journal: ApJ
@@ 1 Introduction
dSph galaxies exhibit a rich diversity in structural, kinematic and chemical parameters (for reviews, see McConnachie, 2012; Simon, 2019; Battaglia & Nipoti, 2022). While the brighter dSph galaxies contain multiple chemo-dynamically distinguishable stellar populations with limited metallicity dispersion (e.g. Tolstoy et al., 2004; Battaglia et al., 2006; Pace et al., 2020), several Milky Way ultra faint galaxies appear to consist of a single stellar population with a relatively large metallicity spread (e.g. Kirby et al., 2013; Frebel et al., 2014; Koch & Rich, 2014; Walker et al., 2016; Kirby et al., 2017), i.e., for those systems, there is no obvious trend with metallicity in the spatial distribution or the kinematics of stars. Developing an understanding of the cause for this diversity is complicated by the multitude of external and internal processes that dSphs are subject to, in particular once accreted onto the Milky Way and subject to its tidal field.
@@
The detailed tidal evolution of dSph galaxies has been shown to depend on the mass profile of the underlying dark matter halo (Peñarrubia et al., 2008, 2010; Errani et al., 2015, 2017), its flattening (Sanders et al., 2018), and the distribution of the stellar binding energies in the dark matter potential (Errani et al., 2022). In addition, the evolution of the underlying dark matter halo will depend on its internal kinematics, and on any potential evolution driven by dark matter self-interactions (evaporation, core collapse, see e.g. Despali et al. 2019; Zeng et al. 2022), or by outflows triggered by stellar feedback (e.g. Pontzen & Governato, 2012; Freundlich et al., 2020; Li et al., 2023). Instead of focusing on the details of each of these processes, some intuition may be gained by categorizing the types of perturbation that may affect a stellar system.
@@
Gravitational perturbations can be broadly categorized by their amplitude and by the time scale on which they act, relative to the corresponding internal (physical and time) scales of the perturbed system. Linear perturbations, in contrast to non-linear ones, are small in amplitude with respect to the gravitational binding energy of the stellar system in question. Adiabatic perturbations, in contrast to impulsive ones, act on time scales that are much longer than the dynamical time scale of the perturbed system. Resonant perturbations have a characteristic time scale that matches an internal time scale of the perturbed system.
@@
The response of stellar (and dark matter) distributions to slow perturbations has been studied using adiabatic invariants: Young (1980) used the combined conservation of angular momentum and radial action to describe the response of a star cluster to a growing black hole; Barnes & White (1984) studied the response of spheroids to a growing disc potential; Blumenthal et al. (1986) and Gnedin et al. (2004) described the adiabatic contraction of a dark matter halo due to accretion of baryons (see also Sellwood & McGaugh 2005 and Katz et al. 2014, who applied the Young 1980 algorithm). More recently, Stücker et al. (2023) used the Young (1980) method to describe the tidal evolution of a dark matter halo in the adiabatic limit. For fast perturbations, the impulsive approximation was applied by Aguilar & White (1985) for the study of the tidal interaction between spherical galaxies; Gnedin & Ostriker (1999) and Gnedin et al. (1999) described the response of stellar systems to tidal shocks, and Taylor & Babul (2001); Drakos et al. (2020); Benson & Du (2022) applied the impulse approximation to model the tidal evolution of dark matter subhalos. For the regime of linear perturbations, Rozier & Errani (2024) used the Kalnajs (1977) matrix method of linear response theory to model the self-consistent response of gravitating systems to a perturbation, showing that adiabatic and impulsive perturbations result in the same final equilibrium state for linear perturbations.
@@
In the present work, we study the conditions necessary for initially chemo-dynamically distinct stellar populations to mix in response to a perturbation. We model the response of mono-energetic stellar tracer populations to linear and non-linear perturbations, both in the adiabatic and in the impulsive regime, making use of controlled $N$-body simulations. This is useful because stellar populations in dSph galaxies may be thought of as a combination of multiple populations of massless tracers with spherical density profiles, mono-energetic distribution functions, and isotropic kinematics. Stellar profiles with complicated energy distributions may be constructed by linear superposition of the mono-energetic results presented here. We discuss the response of mono-energetic stellar tracers to perturbations in dark matter halos with centrally-divergent density cusps, and central constant-density cores.
@@
The paper is structured as follows. In Sec. 2, we introduce the numerical setup used in the present study, and derive a family of stellar density profiles with mono-energetic distribution functions, embedded in cuspy and cored dark matter halos. In Sec. 3, we consider the case of adiabatic perturbations, and show that initially energetically separated stellar populations remain separated. In Sec. 4, we turn our attention to impulsive perturbations, where stellar populations may mix energetically and develop radially-biased kinematics. For stellar tracers embedded in cored halos, we show that inefficient phase mixing results in long-lived oscillations in response to an impulsive perturbation. Finally, in Sec. 5, we summarize our main findings.
@@ 2.1 Distribution function for spherical stellar systems
Consider a star on an orbit in a spherical dark matter potential. As an example, we choose a Hernquist (1990) potential, which in the Dehnen (1993) parametrization corresponds to the case of $\gamma=1$, with density profile
@@
In Eq. 5 we denote by $N_{\star}$ the total number of stars, and $r_{\star}$ is a scale radius. We refer to the radius which encloses half of the stars as the half light radius, which for an exponential profile equals $r_{\mathrm{h}}\approx 2.67\,r_{\star}$. The number density profile is shown in the top panel of Fig. 1 (see also Fig. 12 in Errani et al. 2024 for a comparison of this profile against observed stellar densities of Milky Way dSphs).
@@
where we have assumed that the stellar number density $\nu_{\star}(r)$ vanishes sufficiently fast towards infinity so that $\left.{\mathrm{d}\nu_{\star}}/{\mathrm{d}\psi}\right|_{\psi=1}=0$. For an exponential stellar tracer that is deeply embedded within the power-law cusp of the Hernquist halo, $r_{\star}\ll a$, the potential may be approximated by $\psi(r)\approx{r}/{a}$ and ${\mathrm{d}^{2}\nu_{\star}}/{\mathrm{d}\psi^{2}}\approx\nu_{\star}(\psi)a^{2}/r_{\star}^{2}$. For these assumptions, the distribution function has a simple analytical form,
@@
peaks at an energy of ^$\hat{\mathcal{E}}_{\star}=(7/2)(r_{\star}/a)$. Note that the functional form of Eq. 10 is identical to the empirically motivated form proposed in Errani et al. (2022) to parameterize stellar energy distributions in LCDM. The differential energy distribution Eq. 10 is shown as a black curve in the bottom panel of Fig. 1.
@@ 2.2 Decomposition into mono-energetic profiles
centred around some energy $\mathcal{E}_{\star}$. We compute the number density profile by integrating the distribution function over all velocities $v$,
@@
$\nu_{\star}(r)/\nu_{\star}(0)=\sqrt{1-\psi(r)/\mathcal{E}_{\star}}~{}.$ (13)
@@ 2.3 Stellar and dark matter profiles considered in this work
which, for $r\ll a$, approaches $-\gamma$, whereas for $r\gg a$, the logarithmic slope of the density profile asymptotically approaches $-4$. For the cuspy ($\gamma=1$) profile, the circular velocity curve $V_{\mathrm{c}}(r)=\sqrt{GM(<r)/r}$ peaks at a radius $r_{\mathrm{mx}}=a$ with circular velocity $V_{\mathrm{mx}}=\sqrt{-\Phi_{0}}/2$, and $\Phi_{0}=-GM/a$. For the cored profile ($\gamma=0$), we have $r_{\mathrm{mx}}=2a$ with a peak circular velocity of $V_{\mathrm{mx}}=\sqrt{-8\Phi_{0}/27}$ and $\Phi_{0}=-GM/(2a)$. For both models, $\mathrm{d}\ln\rho/\mathrm{d}\ln r=-2$ at a radius of $2r_{\mathrm{mx}}$.
@@ 3 Adiabatic perturbations
Figure 5: Initial ( ) and final ( ) equilibrium configurations of stellar tracer populations in the numerical experiments with an adiabatically evolving dark matter halo. Shown are half-light radii $r_{\mathrm{h}}$ and average velocity dispersions $\langle\sigma^{2}\rangle^{1/2}$ of four stellar tracers with mono-energetic initial conditions, embedded in a cuspy ($\gamma=1$ in Eq. 1, left panel) and a cored ($\gamma=0$, right panel) dark matter halo. Stellar tracers with initial half-light radii $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$ are shown in dark blue , light blue, orange and red, respectively. All initial conditions lie along the thick gray curve labeled $M/M_{0}=1$. In each simulation, the mass sourcing the underlying dark matter potential is lowered slowly by $-0.2\,\mathrm{dex}$, $-0.4\,\mathrm{dex}$, $\dots$, $-2\,\mathrm{dex}$. Solid gray curves are computed from Eq. 30, assuming mono-energetic distribution functions with isotropic kinematics. All equilibrium configurations fall along lines of $r_{\mathrm{h}}\langle\sigma^{2}\rangle^{1/2}=\text{const}$.
@@ 3.1 Adiabatic evolution of size and kinematics
We will now turn our attention to the evolution of stellar populations subject to slow changes to the underlying potential. Specifically, we study how the four mono-energetic stellar distributions introduced in Sec. 2.3 respond to slow changes to the (dark matter) mass sourcing the underlying potential. We study the evolution separately for stars embedded in a cuspy and a cored dark matter halo. We slowly decrease the total mass $M$ of each dark matter halo, while keeping its radius of peak circular velocity $r_{\mathrm{mx}}$ unaltered (see Eq. 1 for definition). To ensure that the change of the underlying potential is perceived as adiabatic by each stellar population, we decrease the mass $M$ exponentially with a decay constant that is at least 10 times longer than the period of a circular orbit at the (final) half-light radius in the new equilibrium. The results of this experiment are shown in Fig. 5 for a cuspy (left panel) and a cored (right panel) dark matter halo. The snapshots shown correspond to the new equilibrium configuration reached after adiabatically lowering the (dark matter) mass $M$ from $M/M_{0}=1$ to a value of $\log_{10}M/M_{0}=0$, $-0.2$, $\dots$, $-2.0$. Snapshots for the stellar models with initial sizes of $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$ are shown in dark blue, light blue, orange, and red, respectively. The gray curves show the average velocity dispersion $\langle\sigma^{2}\rangle^{1/2}$ expected for a mono-energetic stellar distribution with half-light radius $r_{\mathrm{h}}$.
@@ 3.3 Adiabatic evolution of the energy distribution
The four stellar populations studied in this work have mono-energetic initial conditions. In Fig. 7, we show histograms of the energy distributions for adiabatic changes to the underlying potential, defining energies with respect to the potential minimum in the initial conditions, $\mathcal{E}^{\mathrm{(IC)}}=1-E/\Phi_{0}^{\mathrm{(IC)}}$. With this definition, $\mathcal{E}^{\mathrm{(IC)}}=1$ corresponds to the bound/unbound limit ($E=0$) for both the initial conditions and the evolved potential. The top panel of Fig. 7 shows the initial energy distributions of four stellar components embedded in a cuspy dark matter halo, with initial half-light radii of $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$ shown in dark blue, light blue, orange, and red, respectively. The subsequent panels show the energy distributions in the evolved potential, where the mass sourcing the potential has been adiabatically lowered to $\log_{10}M/M_{0}=-0.2$, $-0.4$, $-1.0$. The evolved potential minimum $\Phi_{0}$ is marked in each panel, and the (inaccessible) range of energies below that minimum is colored gray. The energy distributions remain narrow throughout their evolution. The energetic separation of the four mono-energetic stellar components decreases as (dark matter) mass is removed, but the stellar distributions remain energetically well separated.
The bottom panel of Fig. 7 shows the final equilibrium energy distributions for the simulation with $\log_{10}M/M_{0}=-1.0$ in units of the evolved potential minimum $\mathcal{E}=1-E/\Phi_{0}$. Here, we note a slight broadening of the energy distribution for the more extended tracers with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/16$ and $1/64$. To gain some insight into the reason for the broadening of the distribution, we consider separately the case of circular and radial orbits within the adiabatically evolving potential. For the case of adiabatically evolving spherical potentials, the invariants (see e.g. Weinberg, 1994) are the angular momentum $J$ (Eq. 27) and the radial action $I$ (the latter drifts at second order with the rate of change of the underlying potential, see Peñarrubia 2013 and Burger et al. 2021 for details),
@@
Figure 7: Equilibrium energy distributions of the four stellar profiles embedded in the cuspy halo shown in the right-hand panel of Fig. 5. Energies are expressed in units of the initial potential minimum, $\mathcal{E}^{\mathrm{(IC)}}=1-E/\Phi_{0}^{\mathrm{(IC)}}$. The top panel shows the initial conditions (IC), while the three subsequent panels correspond to snapshots where the mass $M$ sourcing the potential has been slowly lowered by -0.2, -0.4 and $-1.0,\mathrm{dex}$ with respect to the initial value $M_{0}$. The value of the evolved potential minimum $\Phi_{0}$ is shown in each panel, and the inaccessible region of energies below $\Phi_{0}$ is shaded in gray. Bottom panel: energy distribution for the model with $\log_{10}M/M_{0}=-1.0$, with energies expressed in units of the evolved potential minimum $\mathcal{E}=1-E/\Phi_{0}$. The slight broadening of the energy distribution is consistent with the distinct evolution of radial and circular orbits (shown as dashed and dotted vertical lines, respectively). The same systematics are observed in the case of the cored halo (not shown).
As an example, we explicitly show this coincidence of evolved energies for the case of stellar orbits deeply within ($\mathcal{E}\ll 1$) in the cuspy potential explicitly. For a circular orbit of radius $r_{\mathrm{c}}$ within the power-law cusp, $\mathcal{E}=(3/2)~{}r_{\mathrm{c}}/a$ and
[…]

### v2

@@ Abstract
We study the response of mono-energetic stellar populations with initially isotropic kinematics to impulsive and adiabatic changes to an underlying dark matter potential. Half-light radii expand and velocity dispersions decrease as enclosed dark matter is removed. The details of this expansion and cooling depend on the time scale on which the underlying potential changes. In the adiabatic regime, the product of half-light radius and average velocity dispersion is conserved. We show that the stellar populations maintain centrally isotropic kinematics throughout their adiabatic evolution, and their densities can be approximated by a family of analytical radial profiles. Metallicity gradients within the galaxy flatten as dark matter is slowly removed. In the case of strong impulsive perturbations, stellar populations develop power-law-like density tails with radially biased kinematics. We show that the distribution of stellar binding energies within the dark matter halo substantially widens after an impulsive perturbation, no matter the sign of the perturbation. This allows initially energetically separated stellar populations to mix, to the extent that previously chemo-dynamically distinct populations may masquerade as a single population with large metallicity and energy spread. Finally, we show that in response to an impulsive perturbation, stellar populations that are deeply embedded in cored dark matter halos undergo a series of damped oscillations before reaching a virialized equilibrium state, driven by inefficient phase mixing in the harmonic potentials of cored halos. This slow return to equilibrium adds substantial systematic uncertainty to dynamical masses estimated from Jeans modeling or the virial theorem.
Cold dark matter (265); Dwarf spheroidal galaxies (420); Galaxy dynamics (591); NN-body simulations (1083); Perturbation methods (1215)
@@ 1 Introduction
dSph galaxies exhibit a rich diversity in structural, kinematic, and chemical parameters (for reviews, see McConnachie, 2012; Simon, 2019; Battaglia & Nipoti, 2022). While the brighter dSph galaxies contain multiple chemo-dynamically distinguishable stellar populations with limited metallicity dispersion (e.g., Tolstoy et al., 2004; Battaglia et al., 2006; Pace et al., 2020), several Milky Way ultra-faint galaxies appear to consist of a single stellar population with a relatively large metallicity spread (e.g., Kirby et al., 2013; Frebel et al., 2014; Koch & Rich, 2014; Walker et al., 2016; Kirby et al., 2017); i.e., for those systems, there is no obvious trend with metallicity in the spatial distribution or the kinematics of stars. Developing an understanding of the cause for this diversity is complicated by the multitude of external and internal processes that dSphs are subject to, in particular once accreted onto the Milky Way and subject to its tidal field.
@@
The detailed tidal evolution of dSph galaxies has been shown to depend on the mass profile of the underlying dark matter halo (Peñarrubia et al., 2008, 2010; Errani et al., 2015, 2017), its flattening (Sanders et al., 2018), and the distribution of the stellar binding energies in the dark matter potential (Errani et al., 2022). In addition, the evolution of the underlying dark matter halo will depend on its internal kinematics, and on any potential evolution driven by dark matter self-interactions (such as evaporation and core collapse, see, e.g., Despali et al. 2019; Zeng et al. 2022), or by outflows triggered by stellar feedback (e.g. Pontzen & Governato, 2012; Freundlich et al., 2020; Li et al., 2023). Instead of focusing on the details of each of these processes, some intuition may be gained by categorizing the types of perturbation that may affect a stellar system.
@@
Gravitational perturbations can be broadly categorized by their amplitude and by the time scale on which they act, relative to the corresponding internal (physical and time) scales of the perturbed system. Linear perturbations, in contrast to nonlinear ones, are small in amplitude with respect to the gravitational binding energy of the stellar system in question. Adiabatic perturbations, in contrast to impulsive ones, act on time scales that are much longer than the dynamical time scale of the perturbed system. Resonant perturbations have a characteristic time scale that matches an internal time scale of the perturbed system.
@@
The response of stellar (and dark matter) distributions to slow perturbations has been studied using adiabatic invariants: Young (1980) used the combined conservation of angular momentum and radial action to describe the response of a star cluster to a growing black hole; Barnes & White (1984) studied the response of spheroids to a growing disc potential; Blumenthal et al. (1986) and Gnedin et al. (2004) described the adiabatic contraction of a dark matter halo due to accretion of baryons (see also Sellwood & McGaugh 2005 and Katz et al. 2014, who applied the Young 1980 algorithm). More recently, Stücker et al. (2023) used the Young (1980) method to describe the tidal evolution of a dark matter halo in the adiabatic limit. For fast perturbations, the impulsive approximation was applied by Aguilar & White (1985) and Banik & van den Bosch (2021a) for the study of the tidal interaction between spherical galaxies; Gnedin & Ostriker (1999) and Gnedin et al. (1999) described the response of stellar systems to tidal shocks, and Taylor & Babul (2001); Drakos et al. (2020); Benson & Du (2022) applied the impulse approximation to model the tidal evolution of dark matter subhalos. For the regime of linear perturbations, Rozier & Errani (2024) used the Kalnajs (1977) matrix method of linear response theory to model the self-consistent response of gravitating systems to a perturbation, showing that adiabatic and impulsive perturbations result in the same final equilibrium state for linear perturbations.
@@
In the present work, we study the conditions necessary for initially chemo-dynamically distinct stellar populations to mix in response to a perturbation. We model the response of mono-energetic stellar tracer populations to linear and nonlinear perturbations, both in the adiabatic and in the impulsive regime, making use of controlled $N$-body simulations. This is useful because stellar populations in dSph galaxies may be thought of as a combination of multiple populations of massless tracers with spherical density profiles, mono-energetic distribution functions, and isotropic kinematics. Stellar profiles with complicated energy distributions may be constructed by linear superposition of the mono-energetic results presented here. We discuss the response of mono-energetic stellar tracers to perturbations in dark matter halos with centrally-divergent density cusps, and central constant-density cores.
@@
The paper is structured as follows. In Sec. 2, we introduce the numerical setup used in the present study, and derive a family of stellar density profiles with mono-energetic distribution functions, embedded in cuspy and cored dark matter halos. In Sec. 3, we consider the case of adiabatic perturbations, and show that initially energetically separated stellar populations remain separated. In Sec. 4, we turn our attention to impulsive perturbations, where stellar populations may mix energetically and develop radially biased kinematics. For stellar tracers embedded in cored halos, we show that inefficient phase mixing results in long-lived oscillations in response to an impulsive perturbation. Finally, in Sec. 5, we summarize our main findings.
@@ 2.1 Distribution function for spherical stellar systems
Consider a star on an orbit in a spherical dark matter potential. As an example, we choose a Hernquist (1990) potential, which in the Dehnen (1993) parameterization corresponds to the case of $\gamma=1$, with density profile
@@
In Eq. 5 we denote by $N_{\star}$ the total number of stars, and $r_{\star}$ is a scale radius. We refer to the radius that encloses half of the stars as the half-light radius, which, for an exponential profile, equals $r_{\mathrm{h}}\approx 2.67\,r_{\star}$. The number density profile is shown in the top panel of Fig. 1 (see also Fig. 12 in Errani et al. 2024 for a comparison of this profile against observed stellar densities of Milky Way dSphs).
@@
where we have assumed that the stellar number density $\nu_{\star}(r)$ vanishes sufficiently fast toward infinity so that $\left.{\mathrm{d}\nu_{\star}}/{\mathrm{d}\psi}\right|_{\psi=1}=0$. For an exponential stellar tracer that is deeply embedded within the power-law cusp of the Hernquist halo, $r_{\star}\ll a$, the potential may be approximated by $\psi(r)\approx{r}/{a}$ and ${\mathrm{d}^{2}\nu_{\star}}/{\mathrm{d}\psi^{2}}\approx\nu_{\star}(\psi)a^{2}/r_{\star}^{2}$. For these assumptions, the distribution function has a simple analytical form,
@@
peaks at an energy of $\hat{\mathcal{E}}_{\star}=(7/2)(r_{\star}/a)$. Note that the functional form of Eq. 10 is identical to the empirically motivated form proposed in Errani et al. (2022) to parameterize stellar energy distributions in LCDM. The differential energy distribution Eq. 10 is shown as a black curve in the bottom panel of Fig. 1.
@@ 2.2 Decomposition into mono-energetic profiles
centred around some energy $\mathcal{E}_{\star}$. We compute the number density profile by integrating the distribution function over all velocities $v$, which for $\psi(r)\leq\mathcal{E}_{\star}$ gives
@@
$\nu_{\star}(r)/\nu_{\star}(0)=\begin{dcases}\sqrt{1-\psi(r)/\mathcal{E}_{\star}}~&\text{if }\psi(r)\leq\mathcal{E}_{\star}\\ 0~&\text{otherwise.}\end{dcases}$ (13)
@@ 2.3 Stellar and dark matter profiles considered in this work
which, for $r\ll a$, approaches $\gamma$, whereas for $r\gg a$, the logarithmic slope of the density profile asymptotically approaches 4-4. For the cuspy ($\gamma=1$) profile, the circular velocity curve $V_{\mathrm{c}}(r)=\sqrt{GM(<r)/r}$ peaks at a radius $r_{\mathrm{mx}}=a$ with circular velocity $V_{\mathrm{mx}}=\sqrt{-\Phi_{0}}/2$, and $\Phi_{0}=-GM/a$. For the cored profile ($\gamma=0$), we have $r_{\mathrm{mx}}=2a$ with a peak circular velocity of $V_{\mathrm{mx}}=\sqrt{-8\Phi_{0}/27}$ and $\Phi_{0}=-GM/(2a)$. For both models, $\mathrm{d}\ln\rho/\mathrm{d}\ln r=-2$ at a radius of $2r_{\mathrm{mx}}$.
@@ 3 Adiabatic perturbations
Figure 5: Initial ( ) and final ( ) equilibrium configurations of stellar tracer populations in the numerical experiments with an adiabatically evolving dark matter halo. Shown are half-light radii $r_{\mathrm{h}}$ and average velocity dispersions $\langle\sigma^{2}\rangle^{1/2}$ of four stellar tracers with mono-energetic initial conditions, embedded in a cuspy ($\gamma=1$ in Eq. 1, left panel) and a cored ($\gamma=0$, right panel) dark matter halo. Stellar tracers with initial half-light radii $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$ are shown in dark blue, light blue, orange and red, respectively. All initial conditions lie along the thick gray curve labeled $M/M_{0}=1$. In each simulation, the mass sourcing the underlying dark matter potential is lowered slowly by $0.2\,\mathrm{dex}$, $0.4\,\mathrm{dex}$, $\dots$, $2\,\mathrm{dex}$. Solid gray curves are computed from Eq. 30, assuming mono-energetic distribution functions with isotropic kinematics. All equilibrium configurations fall along lines of $r_{\mathrm{h}}\langle\sigma^{2}\rangle^{1/2}=\text{const}$.
@@ 3.1 Adiabatic evolution of size and kinematics
We will now turn our attention to the evolution of stellar populations subject to slow changes to the underlying potential. Specifically, we study how the four mono-energetic stellar distributions introduced in Sec. 2.3 respond to slow changes to the (dark matter) mass sourcing the underlying potential. We study the evolution separately for stars embedded in a cuspy and a cored dark matter halo. We slowly decrease the total mass $M$ of each dark matter halo, while keeping its radius of peak circular velocity $r_{\mathrm{mx}}$ unaltered (see Eq. 1 for definition). To ensure that the change of the underlying potential is perceived as adiabatic by each stellar population, we decrease the mass $M$ exponentially with a decay constant that is at least 10 times longer than the period of a circular orbit at the (final) half-light radius in the new equilibrium. The results of this experiment are shown in Fig. 5 for a cuspy (left panel) and a cored (right panel) dark matter halo. The snapshots shown correspond to the new equilibrium configuration reached after adiabatically lowering the (dark matter) mass $M$ from $M/M_{0}=1$ to a value of $\log_{10}M/M_{0}=0$, 0.2-0.2, $\dots$, 2.0-2.0. Snapshots for the stellar models with initial sizes of $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$ are shown in dark blue, light blue, orange, and red, respectively. The gray curves show the average velocity dispersion $\langle\sigma^{2}\rangle^{1/2}$ expected for a mono-energetic stellar distribution with half-light radius $r_{\mathrm{h}}$.
@@ 3.3 Adiabatic evolution of the energy distribution
The four stellar populations studied in this work have mono-energetic initial conditions. In Fig. 7, we show histograms of the energy distributions for adiabatic changes to the underlying potential, defining energies with respect to the potential minimum in the initial conditions, $\mathcal{E}^{\mathrm{(IC)}}=1-E/\Phi_{0}^{\mathrm{(IC)}}$. With this definition, $\mathcal{E}^{\mathrm{(IC)}}=1$ corresponds to the bound/unbound limit ($E=0$) for both the initial conditions and the evolved potential. The top panel of Fig. 7 shows the initial energy distributions of four stellar components embedded in a cuspy dark matter halo, with initial half-light radii of $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$ shown in dark blue, light blue, orange, and red, respectively. The subsequent panels show the energy distributions in the evolved potential, where the mass sourcing the potential has been adiabatically lowered to $\log_{10}M/M_{0}=-0.2$, 0.4-0.4, 1.0-1.0. The evolved potential minimum $\Phi_{0}$ is marked in each panel, and the (inaccessible) range of energies below that minimum is colored gray. The energy distributions remain narrow throughout their evolution. The energetic separation of the four mono-energetic stellar components decreases as (dark matter) mass is removed, but the stellar distributions remain energetically well separated.
The bottom panel of Fig. 7 shows the final equilibrium energy distributions for the simulation with $\log_{10}M/M_{0}=-1.0$ in units of the evolved potential minimum $\mathcal{E}=1-E/\Phi_{0}$. Here, we note a slight broadening of the energy distribution for the more extended tracers with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/16$ and $1/64$. To gain some insight into the reason for the broadening of the distribution, we consider separately the case of circular and radial orbits within the adiabatically evolving potential. For the case of adiabatically evolving spherical potentials, the invariants (see e.g. Weinberg, 1994a) are the angular momentum $J$ (Eq. 27) and the radial action $I$ (the latter drifts at second order with the rate of change of the underlying potential, see Peñarrubia 2013 and Burger et al. 2021 for details),
@@
Figure 7: Equilibrium energy distributions of the four stellar profiles embedded in the cuspy halo shown in the right-hand panel of Fig. 5. Energies are expressed in units of the initial potential minimum, $\mathcal{E}^{\mathrm{(IC)}}=1-E/\Phi_{0}^{\mathrm{(IC)}}$. The top panel shows the initial conditions (IC), while the three subsequent panels correspond to snapshots where the mass $M$ sourcing the potential has been slowly lowered by -0.2, -0.4 and $1.0,\mathrm{dex}$ with respect to the initial value $M_{0}$. The value of the evolved potential minimum $\Phi_{0}$ is shown in each panel, and the inaccessible region of energies below $\Phi_{0}$ is shaded in gray. Bottom panel: energy distribution for the model with $\log_{10}M/M_{0}=-1.0$, with energies expressed in units of the evolved potential minimum $\mathcal{E}=1-E/\Phi_{0}$. The slight broadening of the energy distribution is caused by the difference in response of radial and circular orbits to the adiabatic perturbation. Dashed and dotted vertical lines are computed from the conservation of radial action (Eq. 35) and angular momentum (Eq. 27), respectively, and accurately delimit the energy distributions measured in the $N$-body snapshots. The same systematics are observed in the case of the cored halo (not shown).
As an example, we explicitly show the coinciding of evolved energies for the case of stellar orbits deeply within ($\mathcal{E}\ll 1$) the cuspy potential. For a circular orbit of radius $r_{\mathrm{c}}$ within the power-law cusp, $\mathcal{E}=(3/2)~r_{\mathrm{c}}/a$ and
$J_{\gamma=1}\stackrel{{\scriptstyle\begin{subarray}{c}\mathcal{E}\ll 1\end{subarray}}}{{\approx}}a\sqrt{-\Phi_{0}}~\left(2\mathcal{E}/3\right)^{3/2}~~~\text{(circular orbit)}~.$ (36)